OpenStudy (anonymous):

I DO NOT understand the algebra that I am doing and would GREATLY appreciate step by step help if anyone would help me. The equations is: Graph the line passing through the given point and having the given slope. Give the slope-intercept form of the equation of the line if possible. (1,-2), m=-3/5 The point-slope form of the equation of a line with slope m passing through the point is the part I really can't figure out. I don't understand how to distribute the properties when fractions are involved. Will anyone please help me???

OpenStudy (anonymous):

Ok, well point slope is just: \[y-k = m(x-h)\] Where m is the slope, and (h.k) is any point on the line.

OpenStudy (amistre64):

given a point P(x,y) and a slope m y-Py = m(x-Px)

OpenStudy (anonymous):

For this problem you would have: \[y-(-2) = -\frac{3}{5}(x-1)\] And to get slope intercept from that you just solve for y. You are having trouble distributing the fractional slope?

OpenStudy (anonymous):

y-(-2)=-3/5(x-1) is where I get stuck because I do not understand the distributing to get to y= for the slope-intercept form of the equation of the line.

OpenStudy (anonymous):

Ok, so to distribute you just multiply x and -1 by -3/5

OpenStudy (amistre64):

i tend to perfer my version of the slope intercept form; y = m(x-Px)+Py which simply skips a few steps of algebra

OpenStudy (anonymous):

Yeah, but the problem is with distributing fractions which will come up in more places than just this.. ;p

OpenStudy (anonymous):

\[y-(-2) = -\frac{3}{5}(x-1)\]\[\implies y -(-2) = -\frac{3}{5}(x) - \frac{3}{5}(-1)\] Right? That's just simple distribution. Is the question how to do the multiplication of a fraction and a non-fraction?

OpenStudy (anonymous):

That is where I am completely stuck!

OpenStudy (anonymous):

Ok well when multiplying a fraction and a whole number just multiply the whole number by the numerator. Then simplify. \[\frac{a}{b}(c) = \frac{a\times c}{b}\]

OpenStudy (anonymous):

For the x you can actually just leave it as \[-\frac{3}{5}x\]

OpenStudy (anonymous):

If I type out the problem that I am working on now, would you mind looking at the steps to see if I am doing them correctly?

OpenStudy (anonymous):

Not at all.

OpenStudy (anonymous):

OK, I started out with (1,2), m=-1/4. After plotting the line, I got the other point which is (5,-3). Now I have to figure out the slope-intercept form of the equation of the line.

OpenStudy (anonymous):

You don't need to get another point if you are given the slope

OpenStudy (anonymous):

Just plug it into the point slope formula.

OpenStudy (anonymous):

Now the book says to use the formula, (y-y1)=m(x-x1)

OpenStudy (anonymous):

Right, m is the slope. y1 is the y part of your point and x1 is the x part of your point. You don't need a second point.

OpenStudy (anonymous):

y-2=-4/3(x+3) and here's where I get totally stuck!

OpenStudy (anonymous):

Would you mind explaining the steps from here?

OpenStudy (anonymous):

That's a different line..

OpenStudy (anonymous):

Which one are we working on

OpenStudy (anonymous):

That is the way the book has it written for the above problem that we have been working on.

OpenStudy (anonymous):

Oops! That was the example. I'm getting mixed up going back and forth.

OpenStudy (anonymous):

"I started out with (1,2), m=-1/4" \[\implies y-2 = -\frac{1}{4}(x-1)\]

OpenStudy (anonymous):

That is the type of form that the problem should be in.

OpenStudy (anonymous):

OK, thank you. This is where I don't understand the multiplication and subtraction to isolate y=.

OpenStudy (anonymous):

Ok so start by distributing the -1/4

OpenStudy (anonymous):

y-2=-1/4x-1/4?

OpenStudy (anonymous):

Close. \[-\frac{1}{4}(-1) = \frac{1}{4}\]

OpenStudy (anonymous):

OK so y-2=-1/4x+1/4?

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

Next would be, y-2-2=-1/4x+1/4-2?

OpenStudy (anonymous):

No, because y-2-2 = y-4

OpenStudy (anonymous):

OK, y-2+2=-1/4x+1/4+2? to isolate y

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

How do I subtract on the right side of the equation?

OpenStudy (anonymous):

You mean add?

OpenStudy (anonymous):

Yes, sorry.

OpenStudy (anonymous):

\[2 = \frac{8}{4}\]\[\implies \frac{1}{4} + 2 = \frac{1}{4} + \frac{8}{4} =\ ?\]

OpenStudy (anonymous):

You need to review addition and multiplication of fractions and whole numbers.

OpenStudy (anonymous):

I think those will help you a lot with this

OpenStudy (anonymous):

Thank you so much for your help.

OpenStudy (anonymous):

Would the y slope-intercept form of the equation of the line be y=-1/4x+8/4 or would I simplify?

OpenStudy (anonymous):

No, that's not right. 2 = 8/4, 2 + 1/4 = 8/4 + 1/4 = 9/4

OpenStudy (anonymous):

y=-1/4x+9/4

OpenStudy (anonymous):

Yes. That's it.

OpenStudy (anonymous):

y=-1/4x+9/4, when I input it, its says that is incorrect.

OpenStudy (anonymous):

Careful with the parens.

OpenStudy (anonymous):

Technically it should be y=(-1/4)x + (9/4)

OpenStudy (anonymous):

Thought the parens on the last term don't matter. If it's still claims that's wrong then you have something wrong with the original problem.

OpenStudy (anonymous):

it says the correct answer is y=-1/4x-7/4

OpenStudy (anonymous):

Then you wrote the problem wrong.. Was it by chance: m = -1/4 point is (1,-2) ?

OpenStudy (anonymous):

That is why I am not understanding the distribution. Every problem is like this and you enter the steps one at a time and it has been correct up to this point by the way we did the problem.

OpenStudy (anonymous):

What we did was correct for the problem you stated. I think you miss-stated the original problem.

OpenStudy (anonymous):

The answer they gave would be correct if the point was (1,-2) not (1,2) as you said originally.

OpenStudy (anonymous):

Oh my gosh, you are right. It was -2.

OpenStudy (anonymous):

So the distribution is fine. You understand ok. You just have to make sure you're solving the right problem ;p

OpenStudy (anonymous):

Lol, sorry. I am getting mixed up because every time I look at the "help me solve this" section, after I look at it, it then changes the problem I have to solve.

OpenStudy (anonymous):

np. Did you want to try another one?

OpenStudy (anonymous):

I want to double check this one since I got the negative wrong but I am going to try to do the whole problem and see if I did it correctly. Here we go....

OpenStudy (anonymous):

(y-(-2))=-1/4(x-1)

OpenStudy (anonymous):

y+2=-1/4x+1/4?

OpenStudy (anonymous):

y+2-2=-1/4x+1/4-2?

OpenStudy (anonymous):

sorry one sec.

OpenStudy (anonymous):

How much have I messed up so far, lol?

OpenStudy (anonymous):

Yes, that's right so far.

OpenStudy (anonymous):

OK, here comes the part that gets me every time!!!

OpenStudy (anonymous):

y=-1/4x+2/4???

OpenStudy (anonymous):

Is 2 more than 1/4?

OpenStudy (anonymous):

Which would you rather have.. 2 pies, or 1/4th of a pie.

OpenStudy (anonymous):

Well 2 pies but I don't know if that is right, should it be simplified to 1/2?

OpenStudy (anonymous):

Hold on. You've made a mistake. I'm just trying to show you how you can tell.

OpenStudy (anonymous):

Do I add 2+1 of the fraction which would make it 3/4 instead of 2/4?

OpenStudy (anonymous):

If 2 is more than 1/4th, then if you have 1/4 and you subtract 2 will the result be positive or negative?

OpenStudy (anonymous):

Negative, I think.

OpenStudy (anonymous):

Right, but your answer was positive.

OpenStudy (anonymous):

2 = 8/4 So you have 1/4 - 2. That's the same as 1/4 - 8/4. What is 1/4 minus 8/4?

OpenStudy (anonymous):

7/4

OpenStudy (anonymous):

I'm sorry that I am terrible at this and I appreciate your patience very much!

OpenStudy (anonymous):

-7/4

OpenStudy (anonymous):

if you have 1 fourth and you take 8 fourths way you will be at -7 fourths.

OpenStudy (anonymous):

So on mine, would it be 1/4-2=-1/4

OpenStudy (anonymous):

No..

OpenStudy (anonymous):

2 is 8 fourths

OpenStudy (anonymous):

\[\frac{2}{1} = \frac{8}{4}\]

OpenStudy (anonymous):

So the problem you're having here is addition and subtraction of fractions.

OpenStudy (anonymous):

You need to study this more.

OpenStudy (anonymous):

If I have 1/4th of a dollar. And you tell me that I owe you two dollars. I can pay you my 1/4 and I'd still owe you 1/4 - 8/4 = -7/4ths of a dollar.

OpenStudy (anonymous):

Watch the video's I have linked. They should help.

OpenStudy (anonymous):

The problem you're having doesn't have anything to do with what you're learning about right now. You are stuck on something you're supposed to have learned already, but it got missed. You have to go back and learn this and everything else will become much much easier.

OpenStudy (anonymous):

I have to go. Good luck, and let me know if you want to work on this more later.